# On annihilator multiplication modules

### v1

@@ Abstract.
An $A$-module $E$ is an annihilator multiplication module if, for each $e\in E$, there is a finitely generated ideal $I$ of $A$ such that $ann(e)=ann(IE)$. In this paper, we investigate fundamental properties of annihilator multiplication modules and employ them as a framework for characterizing significant classes of rings and modules, including torsion-free modules, multiplication modules, injective modules, and principal ideal von Neumann regular rings. In addition, we establish that, for such modules, the equality $Ass_{A}(E)=Ass(A)$ holds, thereby providing a precise connection between module-theoretic and ring-theoretic prime structures.
@@ 1. Introduction
In this paper, we focus only on commutative rings with a nonzero identity and nonzero unital modules. In particular, $A$ will always represent such a ring, and $E$ will represent such an $A$-module. The concept of the multiplication module was first introduced and studied by Barnard in [9]. An $A$-module $E$ is called a multiplication module if every submodule $V$ of $E$ has the form $IE$ for some ideal $I$ of $A$. Subsequently, El-Bast and Smith, in their influential paper [15], explored further properties and characterizations of multiplication modules. The notion of multiplication modules provides a way to associate a submodule $V$ of $E$ with an ideal $(V:_{A}E)$ of $A$, where $(V:_{A}E)=ann_{A}(V/E)$. The algebraic properties of $V$ and $(V:_{A}E)$ are frequently compatible. Many authors have since used multiplication modules as a bridge connecting module theory with ring theory. See, for example, [1], [2], [4], [8], [13], [19]-[22], [26], [27], [30]. A proper submodule $V$ of $E$ is called a prime submodule if for each $a\in A$ and $e\in E$, the condition $ae\in V$ implies $a\in(V:_{A}E)$ or $e\in V$ [29]. The set of all prime submodules of $E$ is denoted by $Spec(E)$. For any submodule $K$ of $E$, the variety of $K$ is defined as $V(K)=\{P\in Spec(E):K\subseteq P\}$. While these varieties generate the Zariski topology in commutative ring theory, they may not always define a topology in general modules. However, if $M$ is a multiplication module, these varieties fulfill the axioms for closed sets of a topology on $Spec(M)$. This is known as the quasi Zariski topology of $M$ [26]. Thus, prime submodules have applications in general topology through the concept of multiplication modules.
@@
Multiplication modules are significant because they encompass several important classes of modules, including simple modules, cyclic modules, and von Neumann regular rings and modules. A ring $A$ (not necessarily commutative) is defined as von Neumann regular if, for each $a$ in $A$, there exists $x$ in $A$ such that $a=axa$ [32]. Von Neumann originally introduced these rings due to their connection with continuous geometry [32]. Their algebraic significance has become increasingly recognized in recent years. In 2018, Jayaram and Tekir extended the concept of von Neumann regular rings to modules by introducing $E$-von Neumann regular and weak idempotent elements. Specifically, an element $a$ in $A$ is $E$-von Neumann regular (or weak idempotent) if $aE=a^{2}E$ (or $a-a^{2}\in ann_{A}(E)$). An $A$-module $E$ is called a von Neumann regular (vn-regular) module if, for each $e$ in $E$, $Ae=aE=a^{2}E$ [17]. Every vn-regular module is also a multiplication module. Furthermore, Jayaram and Tekir demonstrated that a finitely generated $A$-module $E$ is vn-regular if and only if, for each $e$ in $E$, there exists a weak idempotent $a$ in $A$ such that $Ae=aE$ [17, Lemma 5]. In functional analysis, Baer *-rings serve as the algebraic counterparts of von Neumann algebras. Every von Neumann algebra is a Baer *-ring because the right annihilator of any subset is generated by a projection. This relationship underscores the structural importance of projections in operator algebras and positions Baer *-rings as an algebraic framework for the study of annihilators and projection lattices. Kaplansky defined a Baer ring as a ring $A$ in which the annihilator $ann(S)=\{a\in A:aS=0\}$ of any subset $S$ of $A$ is generated by an idempotent $b$ in $A$ [23]. Subsequently, Kist refined this definition by considering $S$ as a singleton, rather than an arbitrary subset. According to Kist [24], a ring $A$ is a Baer ring (also referred to as p.q. Baer or P.P. ring) if, for every $a$ in $A$, there exists an idempotent $b$ in $A$ such that $ann(a)=bA$. The present paper adopts Kist’s definition of a Baer ring.
@@
Let $V$ be a submodule of $E$ and $K$ a nonempty subset of $E$. The residual of $V$ by $K$ is denoted by $(V:_{A}K)=\left\{a\in A:aK\subseteq V\right\}$. When $V=(0)$ and $K$ is a submodule of $E$, we prefer $ann(K)$ to denote $((0):K)$. In particular, for each $e\in E$, we use $ann(e)$ instead of $ann(Ae)$. An $A$-module $E$ is said to be a torsion free if $ann(e)=0$ for every $0\neq e\in E$ [29]. $E$ is said to be a faithful module if $ann(E)=0$. It is clear that all torsion-free modules are faithful, and the converse is not true in general. For instance, $\mathbb{Z}$-module $\mathbb{Z}\oplus\mathbb{Z}_{n}$, where $n\geq 2$ is a faithful module which is not torsion-free. In a recent study, Jayaram, Tekir, and Koç introduced Baer modules, a new generalization of von Neumann regular modules, and extended various properties of Baer rings to these modules. Recall from [18] that an $A$-module $E$ is said to be a Baer module if for each $e\in E$ there exists a weak idempotent $a\in A$ such that $ann(e)E=aE$. Also, they defined an $A$-module $E$ to be an annihilator multiplication if for each $e\in E$, there exists a finitely generated ideal $I$ of $A$ such that $ann(e)=ann(IE)$. Among the results in their paper, the authors showed that every finitely generated von Neumann regular module is a Baer module [18, Proposition 2.8], and also they proved that every finitely generated Baer module is also an annihilator multiplication module [18, Lemma 2.1]. Furthermore, [18] established that annihilator multiplication modules provide a link between the theory of Baer modules and Baer rings. Specifically, a finitely generated $A$-module $E$ is a Baer module if and only if $E$ is an annihilator multiplication module and $A/ann(E)$ is a Baer ring [18, Theorem 2.14]. In this article, we aim to investigate additional algebraic properties of annihilator multiplication modules. In Section 2, we analyze the relationships between annihilator multiplication modules and other classical modules, including multiplication modules, von Neumann regular modules, finitely generated Baer modules, torsion-free modules, and simple modules (see Example 1 and Example 2). We also examine the stability of annihilator multiplication modules under direct products, direct sums, quotients, homomorphisms, factor modules, and polynomial modules (see Proposition 2, Proposition 5, Proposition 3, Proposition 4, Corollary 2, Proposition 6). In addition, we study 1-absorbing prime ideals and classical 1-absorbing prime submodules in annihilator multiplication modules (see Proposition 7 and Proposition 8). Section 3 is devoted to the characterization of important classes of rings and modules, such as torsion-free modules, multiplication modules, injective modules, and principal ideal von Neumann regular rings, using annihilator multiplication modules (see Theorem 1, Proposition 9, Proposition 10, Theorem 2). Finally, we prove that $Ass_{A}(E)=Ass(A)$ in annihilator multiplication modules, where $Ass_{A}(E)$ and $Ass(A)$ denote the sets of associated primes of $E$ and $A$, respectively (see Theorem 3).
@@ 2. Basic properties of annihilator multiplication modules
This section outlines the fundamental properties of annihilator multiplication modules. The discussion begins with the formal definition of this concept.
@@
An $A$-module $E$ is said to be an annihilator multiplication module if for each $e\in E$, there exists a finitely generated ideal $I$ of $A$ such that $ann(e)=ann(IE)$.
@@
(i):(i):\ Let $E$ be a multiplication module and choose $e\in E$. Then we have $Ae=(Ae:E)E$. Then we can write $e=a_{1}e_{1}+a_{2}e_{2}+\cdots+a_{n}e_{n}$ for some $e_{1},e_{2},\ldots,e_{n}\in(Ae:E)$. Now, put $I={\textstyle\sum\limits_{i=1}^{n}}Aa_{i}$. Let $v=be\in Ae$. Then we have $v=be=a_{1}(be_{1})+a_{2}(be_{2})+\cdots+a_{n}(be_{n})\in IE$ which implies that $Ae\subseteq IE\subseteq(Ae:E)E\subseteq Ae$. Thus, we have $Ae=IE$ for some finitely generated ideal $I$ of $A$, and thus we get $ann(e)=ann(IE)$, that is, $E$ is an annihilator multiplication module. The rest follows from the fact that every von Neumann regular module is a multiplication module.
(ii):(ii):\ Follows from [18, Lemma 2.11].
(iii):(iii):\ Follows from [18, Example 2.10 (i)].
@@
Consider $\mathbb{Z}$-module $E=\mathbb{Z}_{2}\oplus\mathbb{Z}_{4}$. Then note that $E$ is not a multiplication module, and also not a reduced module since $2^{2}(\overline{0},\overline{1})=(\overline{0},\overline{0})$ and $2(\overline{0},\overline{1})\neq(\overline{0},\overline{0})$. As $E$ is finitely generated, by [18, Proposition 2.7 and Proposition 2.8], $E$ is neither a Baer module nor a von Neumann regular module. On the other hand, $E$ is not a torsion free module and not a simple module. Let $e=(\overline{x},\overline{y})\in E$. Then it is clear that $ann(e)=4\mathbb{Z}$, $2\mathbb{Z}$ or $\mathbb{Z}$. Now we have 3 cases. Case 1: If $ann(e)=4\mathbb{Z}$, then put $I=\mathbb{Z}$ and note that $ann(e)=ann(IE)$. Case 2: If $ann(e)=2\mathbb{Z}$, then put $I=2\mathbb{Z}$ and note that $ann(e)=ann(IE)$ Case 3: If $ann(e)=\mathbb{Z}$, then put $I=0\mathbb{Z}$ and note that $ann(e)=ann(IE)$. By above all cases, $E$ is an annihilator multiplication module.
Let $E$ be a multiplication module and $V$ a finitely generated submodule of $E$. According to [14, Lemma 3.5], there exists a finitely generated ideal $I$ of $A$ such that $V=IE$. The following result provides an analogous statement for annihilator multiplication modules.
@@
(ii) If $V$ is a submodule of $E$ (not necessarily a finitely generated), then there exists an ideal $I$ of $A$ such that $ann(V)=ann(IW)$.
@@
Let $E$ be an $A$-module and $T\subseteq R$ a multiplicatively closed set. Then we denote the quotient module at $T$ by $T^{-1}E=\left\{\frac{e}{t}:e\in E,t\in T\right\}$ over the quotient ring $T^{-1}A$ [29].
@@
Let $p$ be a prime number and consider $\mathbb{Z}$-module $E=\mathbb{Z}_{p}\oplus\mathbb{Q}.\$First note that $E$ is not a finitely generated module. Choose $e=(\overline{1},0)$ and note that $ann(e)=p\mathbb{Z}$. Since $\mathbb{Z}$ is a PID, every ideal II\ has the form $I=n\mathbb{Z}$ for some $n\in\mathbb{Z}.\$Then we have two cases. Case 1: If $n=0,$ then $ann(IE)=\mathbb{Z}\neq p\mathbb{Z}.\$Case 2: If $n\neq 0,\$then $ann(IE)=(0)\neq p\mathbb{Z}$. This shows that $E$ is not an annihilator multiplication module. Let $T=reg(\mathbb{Z})=\mathbb{Z}-\left\{0\right\}.\$Then note that $T^{-1}E$ is a vector space over the field $T^{-1}\mathbb{Z}=\mathbb{Q}.\$Since every vector space is torsion free, by Example 1, $T^{-1}E$ is an annihilator multiplication module.
@@
(ii):(ii):\ Let $e^{\prime}\in E^{\prime}.\$If $e^{\prime}=0$, then we have $ann(e^{\prime})=A=ann((0)E)$, we are done. So assume that $e^{\prime}$ is nonzero. Since $\phi$ is surjective, we can write $e^{\prime}=\phi(e)\$for some $e\in E.\$As EE\ is an annihilator multiplication module, there exists a finitely generated ideal II\ of AA\ such that ann(e)=ann(IE).ann(e)=ann(IE).\ Since $ann(E)=ann(E^{\prime}),\$we conclude that $ann(IE^{\prime})=ann(IE)=ann(e)\subseteq ann(\phi(e)).\$Now, let $a\in ann(\phi(e)).\$Then we have $a\phi(e)=\phi(ae)=0$ which implies thath $ae\in Ker(\phi).\$Since $Ker(\phi)\$is a prime submodule, we have either $a\in(Ker(\phi):E)$ or $e\in Ker(\phi).\$Case 1: Let $a\in(Ker(\phi):E).\$Then we have $aE\subseteq Ker(\phi),\$that is, $\phi(aE)=a\phi(E)=aE^{\prime}=0$ which implies that $a\in ann(E^{\prime})\subseteq ann(IE^{\prime}).\$Case 2: Let $e\in Ker(\phi),\$that is, $\phi(e)=e^{\prime}=0$ which is a contradiction. Thus, we conclude that $ann(\phi(e))\subseteq ann(IE^{\prime}),\$that is, $an(e^{\prime})=ann(IE^{\prime})$ which completes the proof. ∎
@@
Consider the $\mathbb{Z}$-homomorphism $\pi:\mathbb{Z}\oplus\mathbb{Q}\rightarrow\mathbb{Z}_{p}\oplus\mathbb{Q}$ defined by $\pi(e,\frac{a}{b})=(\overline{e},\frac{a}{b}),\$where pp\ is a prime number. It is clear that $ann(\mathbb{Z}\oplus\mathbb{Q})=(0)=\mathbb{Z}_{p}\oplus\mathbb{Q}.$ Then note that $\pi$ is surjective and $Ker(\pi)=p\mathbb{Z}\oplus(0)$ is not a prime submodule of $\mathbb{Z}\oplus\mathbb{Q}$. On the other hand, $\mathbb{Z}\oplus\mathbb{Q}$ is a torsion free module, by Example 1, so is an annihilator multiplication module. However, by Example 2, $\mathbb{Z}_{p}\oplus\mathbb{Q}$ is not an annihilator multiplication module.
According to [3], a nonzero submodule $V$ of $E$ is defined as an essential submodule, also referred to as a large submodule, if for any submodule $K$ of $E$, the condition $K\cap V=(0)$ implies $K=(0)$. Furthermore, a submodule $V$ of $E$ is termed a pure submodule if $IE\cap V=IV$ for every ideal $I$ of AA\ [3].
@@
(iii):(iii):\ It is enough to show that ann(V)=ann(E).ann(V)=ann(E).\ Let $a\in ann(V).\$Then we have aV=(0).aV=(0).\ Since VV\ is a pure submodule, we conclude that $aV=aE\cap V=(0).\$AsV\ V\ is an essential submodule, we conclude that $aE=(0)$ which implies that $ann(V)\subseteq ann(E).\$Since the reverse inclusion is always hold, we have the equality $ann(V)=ann(E).$

### v2

@@ Abstract.
An $A$-module $E$ is said to be an annihilator multiplication module if for each $e\in E$, there exists a finitely generated ideal $I$ of $A$ such that $ann(e)=ann(IE)$. This class of modules is quite large, as it contains multiplication modules, von Neumann regular modules, finitely generated Baer modules, torsion-free modules, and simple modules. This article provides a comprehensive investigation into the algebraic properties of annihilator multiplication modules, and establishes new characterizations for several important classes of rings/modules, including multiplication modules, torsion-free modules, simple modules, uniserial modules, injective modules and Noetherian von Neumann regular rings. Furthermore, we present a construction method using trivial extensions to produce annihilator multiplication rings that are not multiplication rings. In addition, we prove that, for such modules, the equality $Ass_{A}(E)=Ass(A)$ holds, thereby providing a precise connection between module-theoretic and ring-theoretic prime structures. Finally, we provide various examples to demonstrate the above equality may fail if the condition of being annihilator multiplication module is omitted.
@@ 1. Introduction
The fundamental interplay between the structural properties of a ring and those of its modules is a cornerstone of commutative algebra. A central theme in this area is identifying classes of modules that mirror the ideal structure of their underlying rings. The concept of multiplication modules, first introduced by Barnard in [13] and further developed by El-Bast and Smith in [20], has long served as a crucial bridge in this context. An $A$-module $E$ is said to be a multiplication module if every submodule $V$ of $E$ is of the form $IE$ for some ideal $I$ of $A$. Particularly, an ideal $I$ of a ring $A$ is called a multiplication ideal if it is a multiplication $A$-module, and also $A$ is said to be a multiplication ring if each of its ideals is a multiplication ideal [34]. In such modules, the algebraic properties of $V$ are frequently compatible with those of the ideal $(V:_{A}E)=ann_{A}(E/V)$, allowing for a rich translation of ring-theoretic results into the language of module theory. For this reason, multiplication modules have become a standard tool for researchers seeking to extend classical ring results to a broader module-theoretic setting (See, for example, [1], [3], [5], [12], [16], [35], [36], [40]). Another factor showing the importance of multiplication modules/rings is that these class of rings/modules contains simple modules, cyclic modules, von Neumann regular rings/modules, and also they are used in the characterization of various significiant rings/modules. For instance, arithmetical rings (a ring $A$ is said to be an arithmetical ring if $A_{\mathfrak{m}}$ is a valuation ring for each maximal ideal $\mathfrak{m}$ of $A$ [34]) are characterized by means of multiplication ideals as follows: a ring $A$ is an arithmetical ring if and only if its each finitely generated ideal is a multiplication ideal (See, [34, Page 151, Exercise 18]). Furthermore, a fundamental structural property of multiplication modules is their significance in spectral theory. Specifically, multiplication modules offer a natural framework for extending the classical Zariski topology, which is traditionally defined on the prime spectrum of a ring, to the spectrum of a module. More precisely, while the collection of varieties defined on the spectrum of a general module typically fails to satisfy the axioms for closed sets, the multiplication module property ensures that these varieties form a well-defined topology, which is known as quasi-Zariski topology [35].
Nevertheless, the concept of a multiplication module is limited by its restrictive defining conditions. As a result, it excludes several important structures in module theory, including certain torsion-free modules, free modules of $rank\geq 2$, vector spaces of $dimension\geq 2$, etc. Also, the multiplication module is sometimes insufficient for establishing essential correspondences between rings and modules, especially for Baer structures, since the classical multiplication module property does not adequately connect Baer modules with their underlying Baer rings. In 2022, Jayaram, Tekir, and Koç introduced annihilator multiplication modules to fully reflect the relationship between Baer modules and Baer rings. By overcoming the restrictions of multiplication modules, annihilator multiplication modules provide an extensive framework including multiplication modules, torsion-free modules, free modules (and thus vector spaces), and finitely generated Baer modules. The fundamental distinction of this approach lies in establishing an identity based on the annihilators of factor modules, rather than defining submodules through a direct equality $V=IE$. In this paper, we aim to deepen the theory by investigating this class as a significant entity in its own right. Beyond examining annihilator multiplication modules under various algebraic structures, we demonstrate how this module class can be utilized to characterize several important modules and rings. Furthermore, we prove that annihilator multiplication modules establish a connection between the associated primes of the module and those of the underlying ring. We now begin by providing some notions and notations regarding rings and modules that will be used in the sequel.
@@
Throughout the paper, we focus only on commutative rings with a nonzero identity and nonzero unital modules. Let $A$ always represents such a ring and $E$ represents such an $A$-module. A ring $A$ (not necessarily commutative) is defined as von Neumann regular if, for each $a$ in $A$, there exists $x$ in $A$ such that $a=axa$ [41]. Von Neumann originally introduced these rings due to their connection with continuous geometry [41]. Their algebraic significance has become increasingly recognized in recent years. In 2018, Jayaram and Tekir extended the concept of von Neumann regular rings to modules by introducing $E$-von Neumann regular and weak idempotent elements. Specifically, an element $a$ in $A$ is said to be an $E$-von Neumann regular (or weak idempotent) if $aE=a^{2}E$ (or $a-a^{2}\in ann_{A}(E)$). An $A$-module $E$ is called a von Neumann regular (for short, vn-regular) module if, for each $e$ in $E$, $Ae=aE$ for some $E$-von Neumann regular element $a\in A$ [24]. It is easy to see that every vn-regular module (vn-regular ring) is also a multiplication module (a multiplication ring). Furthermore, Jayaram and Tekir demonstrated that a finitely generated $A$-module $E$ is vn-regular if and only if, for each $e$ in $E$, there exists a weak idempotent $a$ in $A$ such that $Ae=aE$ [24, Lemma 5]. In functional analysis, Baer *-rings serve as the algebraic counterparts of von Neumann algebras. Every von Neumann algebra is a Baer *-ring because the right annihilator of any subset is generated by a projection. This relationship underscores the structural importance of projections in operator algebras and positions Baer *-rings as an algebraic framework for the study of annihilators and projection lattices. Kaplansky defined a Baer ring as a ring $A$ in which the annihilator $ann(S)=\{a\in A:aS=0\}$ of any subset $S$ of $A$ is generated by an idempotent $b$ in $A$ [29]. Subsequently, Kist refined this definition by considering $S$ as a singleton, rather than an arbitrary subset. According to Kist [32], a ring $A$ is a Baer ring (also referred to as p.q. Baer or P.P. ring) if, for every $a$ in $A$, there exists an idempotent $b$ in $A$ such that $ann(a)=bA$. The present paper adopts Kist’s definition of a Baer ring. Let $V$ be a submodule of $E$ and $K$ a nonempty subset of $E$. The residual of $V$ by $K$ is denoted by $(V:_{A}K)=\left\{a\in A:aK\subseteq V\right\}$. When $V=(0)$ and $K$ is a submodule of $E$, we prefer $ann(K)$ to denote $((0):K)$. In particular, for each $e\in E$, we use $ann(e)$ instead of $ann(Ae)$. An $A$-module $E$ is said to be a torsion-free if $ann(e)=0$ for every $0\neq e\in E$ [38]. $E$ is said to be a faithful module if $ann(E)=0$. It is clear that all torsion-free modules are faithful, and the converse is not true in general. For instance, $\mathbb{Z}$-module $\mathbb{Z}\oplus\mathbb{Z}_{n}$, where $n\geq 2$ is a faithful module which is not torsion-free. Recall from [25] that an $A$-module $E$ is said to be a Baer module if for each $e\in E$ there exists a weak idempotent $a\in A$ such that $ann(e)E=aE$. Also, they defined an $A$-module $E$ to be an annihilator multiplication if for each $e\in E$, there exists a finitely generated ideal $I$ of $A$ such that $ann(e)=ann(IE)$. The authors in [25, Proposition 2.8] showed that every finitely generated von Neumann regular module is a Baer module, and also they proved that every finitely generated Baer module is also an annihilator multiplication module (See, [25, Lemma 2.1]). Furthermore, [25] established that annihilator multiplication modules provide a link between the theory of Baer modules and Baer rings. Specifically, a finitely generated $A$-module $E$ is a Baer module if and only if $E$ is an annihilator multiplication module and $A/ann(E)$ is a Baer ring [25, Theorem 2.14].
@@
In this article, we aim to investigate the comprehensive algebraic properties of annihilator multiplication modules. In Section 2, we analyze the relationships between annihilator multiplication modules and other classical modules, including multiplication modules, von Neumann regular modules, finitely generated Baer modules, torsion-free modules, and simple modules (see Example 1 and Example 2). We also examine the stability of annihilator multiplication modules under direct products, in direct sums, under localizations, under homomorphisms, in factor modules, and polynomial modules (see Proposition 2, Proposition 5, Proposition 3, Proposition 4, Corollary 2, Proposition 6). For an ideal $I$ of $A$, we say that $I$ is an annihilator multiplication ideal if it is an annihilator multiplication $A$-module. Also, we define a ring $A$ as an annihilator multiplication ring if its each ideal is an annihilator multiplication ideal. We investigate certain annihilator multiplication ideals and multiplication ideals in trivial ring extension (See, Theorem 1, Theorem 2 and Theorem 3). In terms of these results, we characterize simple modules, and also we provide a construction method using trivial ring extensions to produce annihilator multiplication rings that are not multiplication rings (See, Theorem 3 and Example 6). In Section 3, we study 1-absorbing prime ideals and classical 1-absorbing prime submodules in annihilator multiplication modules (see Proposition 9 and Proposition 8). Also, we characterize some significiant classes of rings and modules, such as torsion-free modules, multiplication modules, injective modules, uniserial modules and Noetherian von Neumann regular rings, using annihilator multiplication modules (See, Theorem 4, Proposition 7, Proposition 10, Corollary 3 and Theorem 5). Finally, we prove that $Ass_{A}(E)=Ass(A)$ in annihilator multiplication modules, where $Ass_{A}(E)$ and $Ass(A)$ denote the sets of associated primes of $E$ and $A$, respectively (see Theorem 6). Also, we provide various examples to demonstrate the above equality may fail if the condition of being annihilator multiplication module is omitted (See, Example 8).
@@ 2. Basic properties of annihilator multiplication modules
An $A$-module $E$ is said to be an annihilator multiplication module if for each $e\in E$, there exists a finitely generated ideal $I$ of $A$ such that $ann(e)=ann(IE)$ [25]. In particular, an ideal $I$ of $A$ is an annihilator multiplication ideal if it is an annihilator multiplication $A$-module. Also, a ring $A$ is said to be an annihilator multiplication ring if its each ideal is an annihilator multiplication ideal.
Now, we start by giving some examples of annihilator multiplication modules.
@@
(i):(i):\ Let $E$ be a multiplication module and choose $e\in E$. Then we have $Ae=(Ae:E)E$. Thus we can write $e=a_{1}e_{1}+a_{2}e_{2}+\cdots+a_{n}e_{n}$ for some $a_{1},a_{2},\ldots,a_{n}\in(Ae:E)$. Now, put $I={\textstyle\sum\limits_{i=1}^{n}}Aa_{i}$. Let $v=be\in Ae$. Then we have $v=be=a_{1}(be_{1})+a_{2}(be_{2})+\cdots+a_{n}(be_{n})\in IE$ which implies that $Ae\subseteq IE\subseteq(Ae:E)E\subseteq Ae$. Thus, we have $Ae=IE$ for some finitely generated ideal $I$ of $A$, and thus we get $ann(e)=ann(IE)$, that is, $E$ is an annihilator multiplication module. The rest follows from the fact that every von Neumann regular module is a multiplication module.
(ii):(ii):\ Follows from [25, Lemma 2.11].
(iii):(iii):\ Follows from [25, Example 2.10 (i)].
@@
Consider $\mathbb{Z}$-module $E=\mathbb{Z}_{2}\oplus\mathbb{Z}_{4}$. Then note that $E$ is not a multiplication module, and also not a reduced module since $2^{2}(\overline{0},\overline{1})=(\overline{0},\overline{0})$ and $2(\overline{0},\overline{1})\neq(\overline{0},\overline{0})$. As $E$ is finitely generated, by [25, Proposition 2.7 and Proposition 2.8], $E$ is neither a Baer module nor a von Neumann regular module. On the other hand, $E$ is not a torsion-free module and not a simple module. Let $e=(\overline{x},\overline{y})\in E$. Then it is clear that $ann(e)=4\mathbb{Z}$, $2\mathbb{Z}$ or $\mathbb{Z}$. Now we have 3 cases. Case 1: If $ann(e)=4\mathbb{Z}$, then put $I=\mathbb{Z}$ and note that $ann(e)=ann(IE)$. Case 2: If $ann(e)=2\mathbb{Z}$, then put $I=2\mathbb{Z}$ and note that $ann(e)=ann(IE)$ Case 3: If $ann(e)=\mathbb{Z}$, then put $I=0\mathbb{Z}$ and note that $ann(e)=ann(IE)$. By above all cases, $E$ is an annihilator multiplication module.
Let $E$ be a multiplication module and $V$ a finitely generated submodule of $E$. According to [19, Lemma 3.5], there exists a finitely generated ideal $I$ of $A$ such that $V=IE$. The following result provides an analogous statement for annihilator multiplication modules.
@@
(ii) If $V$ is a submodule of $E$ (not necessarily a finitely generated), then there exists an ideal $I$ of $A$ such that $ann(V)=ann(IE)$.
@@
Let $E$ be an $A$-module and $T\subseteq A$ a multiplicatively closed set. Then we denote the quotient module at $T$ by $T^{-1}E=\left\{\frac{e}{t}:e\in E,t\in T\right\}$ over the quotient ring $T^{-1}A$ [38].
@@
Let $p$ be a prime number and consider $\mathbb{Z}$-module $E=\mathbb{Z}_{p}\oplus\mathbb{Z}.\$First note that $E$ is a finitely generated module. Choose $e=(\overline{1},0)$ and note that $ann(e)=p\mathbb{Z}$. Since $\mathbb{Z}$ is a PID, every ideal II\ has the form $I=n\mathbb{Z}$ for some $n\in\mathbb{Z}.\$Then we have two cases. Case 1: If $n=0,$ then $ann(IE)=\mathbb{Z}\neq p\mathbb{Z}.\$Case 2: If $n\neq 0,\$then $ann(IE)=(0)\neq p\mathbb{Z}$. This shows that $E$ is not an annihilator multiplication module. Let $T=reg(\mathbb{Z})=\mathbb{Z}-\left\{0\right\}.\$Then note that $T^{-1}E$ is a vector space over the field $T^{-1}\mathbb{Z}=\mathbb{Q}.\$Since every vector space is torsion-free, by Example 1, $T^{-1}E$ is an annihilator multiplication module.
@@
(ii):(ii):\ Let $e^{\prime}\in E^{\prime}.\$If $e^{\prime}=0$, then we have $ann(e^{\prime})=A=ann((0)E^{\prime})$, we are done. So assume that $e^{\prime}$ is nonzero. Since $\phi$ is surjective, we can write $e^{\prime}=\phi(e)\$for some $e\in E.\$As EE\ is an annihilator multiplication module, there exists a finitely generated ideal II\ of AA\ such that ann(e)=ann(IE).ann(e)=ann(IE).\ Since $ann(E)=ann(E^{\prime}),\$we conclude that $ann(IE^{\prime})=ann(IE)=ann(e)\subseteq ann(\phi(e)).\$Now, let $a\in ann(\phi(e)).\$Then we have $a\phi(e)=\phi(ae)=0$ which implies that $ae\in Ker(\phi).\$Since $Ker(\phi)\$is a prime submodule, we have either $a\in(Ker(\phi):E)$ or $e\in Ker(\phi).\$Case 1: Let $a\in(Ker(\phi):E).\$Then we have $aE\subseteq Ker(\phi),\$that is, $\phi(aE)=a\phi(E)=aE^{\prime}=0$ which implies that $a\in ann(E^{\prime})\subseteq ann(IE^{\prime}).\$Case 2: Let $e\in Ker(\phi),\$that is, $\phi(e)=e^{\prime}=0$ which is a contradiction. Thus, we conclude that $ann(\phi(e))\subseteq ann(IE^{\prime}),\$that is, $ann(e^{\prime})=ann(IE^{\prime})$ which completes the proof. ∎
@@
Consider the $\mathbb{Z}$-homomorphism $\pi:\mathbb{Z}\oplus\mathbb{Z}\rightarrow\mathbb{Z}_{p}\oplus\mathbb{Z}$ defined by $\pi(m,n)=(\overline{m},n),\$where pp\ is a prime number. It is clear that $ann(\mathbb{Z}\oplus\mathbb{Z})=(0)=ann(\mathbb{Z}_{p}\oplus\mathbb{Z})$. Then note that $\pi$ is surjective and $Ker(\pi)=p\mathbb{Z}\oplus(0)$ is not a prime submodule of $\mathbb{Z}\oplus\mathbb{Z}$. On the other hand, $\mathbb{Z}\oplus\mathbb{Z}$ is a torsion-free module, by Example 1, so is an annihilator multiplication module. However, by Example 2, $\mathbb{Z}_{p}\oplus\mathbb{Z}$ is not an annihilator multiplication module.
According to [7], a nonzero submodule $V$ of $E$ is defined as an essential submodule, also referred to as a large submodule, if for any submodule $K$ of $E$, the condition $K\cap V=(0)$ implies $K=(0)$. Furthermore, a submodule $V$ of $E$ is called a pure submodule if $IE\cap V=IV$ for every ideal $I$ of AA\ [7].
@@
(iii):(iii):\ It is enough to show that ann(V)=ann(E).ann(V)=ann(E).\ Let $a\in ann(V).\$Then we have aV=(0).aV=(0).\ Since VV\ is a pure submodule, we conclude that $aV=aE\cap V=(0).\$AsV\ V\ is an essential submodule, we conclude that $aE=(0)$ which implies that $ann(V)\subseteq ann(E).\$Since the reverse inclusion always holds, we have the equality $ann(V)=ann(E).$